OpenStudy (anonymous):

Can you help me with this: Match the oxidation numbers of nitrogen for the following: 31. NO2 32. No 33. N2O4 34. NO3 35. N3O6 A: +4 B: +2 C: +6

OpenStudy (anonymous):

a.+4 b.+2 c.+6 1.NO2=A 2.NO=B 3.N2O4=A 4.NO3=C 5.N3O6=A

OpenStudy (anonymous):

Just remember that one atom of oxygen has oxidation no = -2 In every case, let x be the oxidation no of one nitrogen For 1)NO2 x + 2(-2) = 0 // 0 is the total charge of the molecule x - 4 = 0 x=4 2)NO x + (-2) = 0 x=2 3.N2O4 2(x) + 4(-2) = 0 2(x) = +8 x = 4 Rest, you can try by yourself for practice @Pamelaa23

OpenStudy (anonymous):

Thank you. So number 4. would be 6 and number 5 would be 2?

OpenStudy (anonymous):

The 4th answer is right but the 5 th answer you told is wrong. Write the equation you formed for the 5th one

OpenStudy (anonymous):

Wait, no. Number 5 is 4?

OpenStudy (anonymous):

Sorry, yes. @Pamelaa23 , number 5 is 4 :)